How much energy would be released if you (when at rest) were to suddenly transform into pure energy? Consider that your mass is approximately 60 kg.

A. 18 x 10^17 J

B. 54x10^17 J

C. 18 x 10 ^9 J

D. 54 x 10^9 J
Answer #1

E = m c^2 = 60 * 3e8 * 3e8 = 54 x 10^17 J
Answer #2

B. 54x10^17 J